For the surface (1/x)+(1/y)+(1/z)= 1 evaluate (partial z)/(partial x) and (partial z)/(partial y) at the point (2,3,6)
1
Expert's answer
2020-10-21T16:32:50-0400
"(1\/x)+(1\/y)+(1\/z)= 1"


Now take partial derivative with respect to x in both sides and using chain rule we get


"-\frac{1}{x^2}-\frac{1}{z^2}\frac{\partial z}{\partial x}=0\\
\implies \frac{\partial z}{\partial x}=-(\frac{z}{x})^2\\
\implies \frac{\partial z}{\partial x}\bigg|_{(2,3,6)}=-(\frac{6}{2})^2=-9"

Now, since "(1\/x)+(1\/y)+(1\/z)= 1" is symmetric in x,y,z thus

"\frac{\partial z}{\partial y}=-(\frac{z}{y})^2\\
\implies \frac{\partial z}{\partial y}\bigg|_{(2,3,6)}=-(\frac{6}{3})^2=-4"